give the equation of the ellipse with intercepts (4,0), (-4,0),(0,2) and (0,-2).
Answer by fractalier(2101)

This ellipse has a center at the origin. That can be seen if you plot the intercepts. Also you can see that the semi-major axis (a) is 4, and the semi-minor axis (b) is 2. So we plug into the standard form of an ellipse to get:
(x^2)/(a^2) + (y^2)/(b^2) = 1
(x^2)/(4^2) + (y^2)/(2^2) = 1
(x^2) / 16 + (y^2) / 4 = 1
and we're done.